# Diffusion Non-Additive Model for Multi-Fidelity Simulations with Tunable Precision

### v1

@@ 7 Conclusion and Discussion
We introduce the Diffusion Non-Additive (DNA) model for multi-fidelity simulations with tunable fidelity parameters, offering a flexible framework for modeling complex dependencies across fidelities. By relaxing the additive structure and allowing for a coupling between the fidelity-controlling parameter and the input space, the DNA model can adapt to a wide range of simulation behaviors. Through comprehensive numerical studies, including synthetic functions and PDE-based simulators, we demonstrate that the DNA model achieves strong predictive performance and reliable uncertainty quantification.
@@
$[\mathbf{K}_{\mathcal{X}}]_{ij}=\exp\left(-\sum_{k=1}^{d}\frac{(x_{ik}-x_{jk})^{2}}{\theta_{j}}\right).$
@@
$[\mathbf{G}_{\delta}]_{ij}=\left(\frac{(t_{i}-t_{j})^{2}}{\theta_{t}}+1\right)^{-\delta},\quad\text{for }2\leq i,j\leq L.$
@@
When $\delta>0$, the function $f_{\delta}(r)=(r+1)^{-\delta}$ is completely monotonic on $[0,\infty)$. By Schoenberg’s theorem (Schoenberg,, 1938), the matrix $\mathbf{G}_{\delta}$ with entries $f_{\delta}((t_{i}-t_{j})^{2}/\theta_{t})$ is strictly positive definite, since the tuning parameters $t_{2},\ldots,t_{L}$ are distinct. Given that $\mathbf{K}_{\mathcal{X}}\succ 0$, the Kronecker product $\mathbf{K}_{\delta}=\mathbf{G}_{\delta}\otimes\mathbf{K}_{\mathcal{X}}$ is also strictly positive definite. Consequently, $\lambda_{\min}(\mathbf{K}_{\delta})>0$. Moreover, since $\mathbf{K}_{\delta}$ is symmetric and finite-dimensional with size $n(L-1)\times n(L-1)$, all its eigenvalues are real and bounded, implying $\lambda_{\max}(\mathbf{K}_{\delta})<\infty$. It follows that $\kappa(\mathbf{K}_{\delta})<\infty$.
@@
$\frac{\partial\mathbf{K}}{\partial\theta_{j}}$ $=-\frac{||x_{j}-x_{j}^{\prime}||^{2}}{\theta_{j}^{2}}\frac{\partial\mathbf{K}}{\partial v},\quad\frac{\partial\mathbf{K}}{\partial\theta_{y}}=-\frac{||y-y^{\prime}||^{2}}{\theta_{y}^{2}}\frac{\partial\mathbf{K}}{\partial v},$
@@
where $u=\left(1+\frac{(t-t^{\prime})^{2}}{\theta_{t}}\right)^{-1}$, $v=\frac{||y-y^{\prime}||^{2}}{\theta_{y}}+\sum_{j=1}^{d}\frac{||x_{j}-x_{j}^{\prime}||^{2}}{\theta_{j}}$, and
@@
where $u=\left(1+\frac{(t-t^{\prime})^{2}}{\theta_{t}}\right)^{-1}$, $v_{j}=\frac{\sqrt{3}||x_{j}-x_{j}^{\prime}||}{\theta_{j}}$, $v_{y}=\frac{\sqrt{3}||y-y^{\prime}||}{\theta_{y}}$, and
@@
where $u=\left(1+\frac{(t-t^{\prime})^{2}}{\theta_{t}}\right)^{-1}$, $v_{j}=\frac{\sqrt{5}||x_{j}-x_{j}^{\prime}||}{\theta_{j}}$, $v_{y}=\frac{\sqrt{5}||y-y^{\prime}||}{\theta_{y}}$, and
@@
×$\quad\quad\quad\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right)$
×$\quad\quad\quad\times\mathbb{E}\left[\exp{\left(-\frac{c_{i}(f_{l-1}(\mathbf{x})-(Y_{-L})_{i})^{2}+c_{k}(f_{l-1}(\mathbf{x})-(Y_{-L})_{k})^{2}}{\theta_{y}}\right)}\Big{|}Y_{N}\right]$ (S3.1)
@@
×$\quad\quad\quad\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right),$
@@
×$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
[…]
@@
[…]
@@
×$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
×$\times\exp\left(-\frac{1}{\left(\frac{(t-t^{\prime})^{2}}{\theta_{t}}+1\right)^{\frac{\beta}{2}}}\frac{\sqrt{5}|x-x^{\prime}|}{\theta}\right).$
@@
where $\psi$ is defined in Section S4, $r_{i}=(\mathbf{K}^{-1}(Y_{-1}-\alpha\mathbf{1}_{N_{-1}}))_{i}$,$c_{i}(t):=c(t,\mathbf{t}_{i})=\left(\frac{(t-\mathbf{t}_{i})^{2}}{\theta_{t}}+1\right)^{-\frac{\beta}{2}}$,$\xi_{i}=\mathbb{E}\left[\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{i};\theta_{y},\nu)\Big{|}Y_{N}\right]$, and $\zeta_{ik}=\mathbb{E}\left[\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{i};\theta_{y},\nu)\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{k};\theta_{y},\nu)\Big{|}Y_{N}\right]$. The closed-form expressions of $\xi_{i}$ and $\zeta_{ik}$ are provided in the following subsections.
@@
×$\times\left[E_{1}^{\prime}\Lambda_{11}\Phi\left(\frac{\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{i}-\frac{\sqrt{3}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right.$
@@
×$\times\left[E_{2}^{\prime}\Lambda_{21}\Phi\left(\frac{-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\frac{\sqrt{3}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right.$
@@
×$\times\left[E_{3}^{\prime}\Lambda_{31}\Phi\left\{\frac{\left(\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{k}-\sqrt{3}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right.\right\}$
@@
×$\times\left[E_{4}^{\prime}\Lambda_{41}\left(\Phi\left\{\frac{(Y_{-L})_{k}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}-\Phi\left\{\frac{(Y_{-L})_{i}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right)\right.$
@@
×$\times\left[E_{5}^{\prime}\Lambda_{51}\Phi\left\{\frac{\left(-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\sqrt{3}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right.$
@@
×$\times\left[E_{1}^{\prime}\Lambda_{11}\Phi\left(\frac{\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{i}-\frac{\sqrt{5}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right.$
@@
×$\times\left[E_{2}^{\prime}\Lambda_{21}\Phi\left(\frac{-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\frac{\sqrt{5}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right.$
@@
×$\times\left[E_{3}^{\prime}\Lambda_{31}\Phi\left\{\frac{\left(\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{k}-\sqrt{5}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right.\right\}$
@@
×$\times\left[E_{4}^{\prime}\Lambda_{41}\left(\Phi\left\{\frac{(Y_{-L})_{k}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}-\Phi\left\{\frac{(Y_{-L})_{i}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right)\right.$
@@
×$\times\left[E_{5}^{\prime}\Lambda_{51}\Phi\left\{\frac{\left(-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\sqrt{5}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right.$
@@
        $\mathbf{k}_{1}(\widetilde{\mathcal{X}}_{1})$ is an $\widetilde{n}_{1}\times n_{1}$ matrix with elements $(\mathbf{k}_{1}(\widetilde{\mathcal{X}}_{1}))_{i,j}=K_{1}((\widetilde{\mathcal{X}}_{1})_{i},\mathbf{x}^{[1]}_{j})$, and $K_{1}(\widetilde{\mathcal{X}}_{1},\widetilde{\mathcal{X}}_{1})$ is an $\widetilde{n}_{1}\times\widetilde{n}_{1}$ matrix with elements $(K_{1}(\widetilde{\mathcal{X}}_{1},\widetilde{\mathcal{X}}_{1}))_{i,j}=K_{1}((\widetilde{\mathcal{X}}_{1})_{i},(\widetilde{\mathcal{X}}_{1})_{j})$.
@@
        $(\displaystyle\mathbf{A}^{(m-1)}_{ij}=K($$((\mathbf{t}^{*}_{-1}\setminus\mathbf{t}_{-1})_{i},(\mathbf{X}^{*}_{-1}\setminus\mathbf{X}_{-1})_{i},(Y^{*(m-1)}_{-L}\setminus Y^{(m-1)}_{-L})_{i}),$
@@
Figure S1: Illustrations of Poisson’s equation (upper panel) and the heat equation (bottom panel) case studies. Each subplots represents subplots with tuning parameter values decreasing from large (left) to zero (right). In each subplot, the black dashed line represents the true function, red dots denote the design points, the blue line indicates the predicted function, and the shaded region depicts the 99% confidence interval.

### v2

@@ 7 Conclusion and Discussion
We introduce the Diffusion Non-Additive (DNA) model for multi-fidelity simulations with tunable fidelity parameters, offering a flexible framework for modeling complex dependencies across fidelities. By relaxing the additive structure and allowing for a coupling between the fidelity-controlling parameter and the input space, the DNA model can adapt to a wide range of simulation behaviors. Through comprehensive numerical studies, including synthetic functions and PDE-based simulators, we demonstrate that the DNA model achieves strong predictive performance and reliable uncertainty quantification. Future work will focus on establishing a rigorous theoretical foundation for the DNA model, including analyses of its asymptotic behavior when $t\rightarrow 0$.
@@
Data Availability Statement The authors confirm that the data supporting the findings of this study are available within its supplementary materials.
@@
$\mathbf{K}_{\mathcal{X}}]_{ij}=\exp\left(-\sum_{k=1}^{d}\frac{(x_{ik}-x_{jk})^{2}}{\theta_{j}}\right).$
@@
[…]
@@
When $\delta>0$, the function $f_{\delta}(r)=(r+1)^{-\delta}$ is completely monotonic on $0,\infty)$. By Schoenberg’s theorem (Schoenberg,, 1938), the matrix $\mathbf{G}_{\delta}$ with entries $f_{\delta}((t_{i}-t_{j})^{2}/\theta_{t})$ is strictly positive definite, since the tuning parameters $t_{2},\ldots,t_{L}$ are distinct. Given that $\mathbf{K}_{\mathcal{X}}\succ 0$, the Kronecker product $\mathbf{K}_{\delta}=\mathbf{G}_{\delta}\otimes\mathbf{K}_{\mathcal{X}}$ is also strictly positive definite. Consequently, $\lambda_{\min}(\mathbf{K}_{\delta})>0$. Moreover, since $\mathbf{K}_{\delta}$ is symmetric and finite-dimensional with size $n(L-1)\times n(L-1)$, all its eigenvalues are real and bounded, implying $\lambda_{\max}(\mathbf{K}_{\delta})<\infty$. It follows that $\kappa(\mathbf{K}_{\delta})<\infty$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\quad\quad\quad\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right)$
$\quad\quad\quad\times\mathbb{E}\left[\exp{\left(-\frac{c_{i}(f_{l-1}(\mathbf{x})-(Y_{-L})_{i})^{2}+c_{k}(f_{l-1}(\mathbf{x})-(Y_{-L})_{k})^{2}}{\theta_{y}}\right)}\Big|Y_{N}\right]$ (S3.1)
@@
$\quad\quad\quad\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right),$
@@
$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
$\frac{1}{\sqrt{2\pi\sigma^{2}}}$ $\int\exp{\left(-Ax^{2}+Bx-C\right)}dx=\frac{1}{\sqrt{2\pi\sigma^{2}}}\int\exp{\left(-A\left(x-\frac{B}{2A}\right)^{2}+\frac{B^{2}}{4A}-C\right)}dx$
$=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp{\left(\frac{B^{2}}{4A}-C\right)}\int\exp{\left(-A\left(x-\frac{B}{2A}\right)^{2}\right)}dx$
@@
$\mathbb{E}$ $\left[\exp{\left(-c_{1}(X-y_{i})^{2}-c_{2}(X-y_{k})^{2}\right)}\Big|Y_{N}\right]=\int\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{\left(-c_{1}(x-y_{i})^{2}-c_{2}(x-y_{k})^{2}\right)}e^{\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right)}dx$
$=\frac{1}{\sqrt{2\pi\sigma^{2}}}\int\exp{\left(-\left(c_{1}+c_{2}+\frac{1}{2\sigma^{2}}\right)x^{2}+\left(2c_{1}y_{i}+2c_{2}y_{k}+\frac{2\mu}{2\sigma^{2}}\right)x-\left(c_{1}y_{i}^{2}+c_{2}y_{k}^{2}+\frac{\mu^{2}}{2\sigma^{2}}\right)\right)}dx$
@@
$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
$\times\exp\left(-\frac{1}{\left(\frac{(t-t^{\prime})^{2}}{\theta_{t}}+1\right)^{\frac{\beta}{2}}}\frac{\sqrt{5}|x-x^{\prime}|}{\theta}\right).$
@@
where $\psi$ is defined in Section S4, $r_{i}=(\mathbf{K}^{-1}(Y_{-1}-\alpha\mathbf{1}_{N_{-1}}))_{i}$$c_{i}(t):=c(t,\mathbf{t}_{i})=\left(\frac{(t-\mathbf{t}_{i})^{2}}{\theta_{t}}+1\right)^{-\frac{\beta}{2}}$$\xi_{i}=\mathbb{E}\left[\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{i};\theta_{y},\nu)\Big|Y_{N}\right]$, and $\zeta_{ik}=\mathbb{E}\left[\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{i};\theta_{y},\nu)\psi(f_{l-1}(\mathbf{x}),(Y_{-L})_{k};\theta_{y},\nu)\Big|Y_{N}\right]$. The closed-form expressions of $\xi_{i}$ and $\zeta_{ik}$ are provided in the following subsections.
@@
$\times\left[E_{1}^{\prime}\Lambda_{11}\Phi\left(\frac{\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{i}-\frac{\sqrt{3}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right$.
@@
$\times\left[E_{2}^{\prime}\Lambda_{21}\Phi\left(\frac{-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\frac{\sqrt{3}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right$.
@@
[…]
@@
$\times\left[E_{4}^{\prime}\Lambda_{41}\left(\Phi\left\{\frac{(Y_{-L})_{k}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}-\Phi\left\{\frac{(Y_{-L})_{i}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right)\right$.
@@
$\times\left[E_{5}^{\prime}\Lambda_{51}\Phi\left\{\frac{\left(-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\sqrt{3}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right$.
@@
$\times\left[E_{1}^{\prime}\Lambda_{11}\Phi\left(\frac{\mu^{*}_{l-1}(\mathbf{x})-(Y_{-L})_{i}-\frac{\sqrt{5}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right$.
@@
$\times\left[E_{2}^{\prime}\Lambda_{21}\Phi\left(\frac{-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\frac{\sqrt{5}c_{i}(t)\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}}{\sigma^{*}_{l-1}(\mathbf{x})}\right)\right$.
@@
[…]
@@
$\times\left[E_{4}^{\prime}\Lambda_{41}\left(\Phi\left\{\frac{(Y_{-L})_{k}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}-\Phi\left\{\frac{(Y_{-L})_{i}-\mu^{*}_{l-1}(\mathbf{x})}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right)\right$.
@@
$\times\left[E_{5}^{\prime}\Lambda_{51}\Phi\left\{\frac{\left(-\mu^{*}_{l-1}(\mathbf{x})+(Y_{-L})_{i}-\sqrt{5}(c_{i}(t)+c_{k}(t))\frac{\sigma^{*2}_{l-1}(\mathbf{x})}{\theta_{y}}\right)}{\sigma^{*}_{l-1}(\mathbf{x})}\right\}\right$.
@@
[…]
@@
        $\mathbf{A}^{(m-1)}_{ij}=K($ $((\mathbf{t}^{*}_{-1}\setminus\mathbf{t}_{-1})_{i},(\mathbf{X}^{*}_{-1}\setminus\mathbf{X}_{-1})_{i},(Y^{*(m-1)}_{-L}\setminus Y^{(m-1)}_{-L})_{i}),$
@@
S7 Supporting figures in Sections 4 and 5
This section provides the supporting tables and figures for Sections 4 and 5, including the comparison results in Section 4 (Figure S1, S2, S3, and S4), and the demonstration of the model predictions with confidence intervals for Poisson’s equation and the heat equation in Section 5 (Figure S5).
Figure S1: RMSEs and CRPSs of four synthetic examples across 100 repetitions with $c=0.75$ and $\gamma=2$.
Figure S2: RMSEs and CRPSs of four synthetic examples across 100 repetitions with $c=\frac{2}{3}$ and $\gamma=1.5$.
Figure S3: RMSEs and CRPSs of four synthetic examples across 100 repetitions with $c=0.55$ and $\gamma=1$.
Figure S4: RMSEs and CRPSs of four synthetic examples across 100 repetitions with $c=0.5$ and $\gamma=1$.
Figure S5: Illustrations of Poisson’s equation (upper panel) and the heat equation (bottom panel) case studies. Each subplot represents subplots with tuning parameter values decreasing from large (left) to zero (right). In each subplot, the black dashed line represents the true function, red dots denote the design points, the blue line indicates the predicted function, and the shaded region depicts the 99% confidence interval.